calculate equilbrium from the following figures c=2000+0.75yd, I=2000, G=4600,X=2550, M=410+0.3Y T=100+0.25Y
1
Expert's answer
2013-06-24T09:45:39-0400
The level of equilibrium national income Ye = C + I + G + (X - IM) , where C= Consumer spending, I = Investment (Gross fixed Capital Formation), G= Government Spending, X= Exports and IM= Imports.
Disposable income ( Yd ) is calculated as the difference between personal income (Y) and personal tax and non tax payments (T).
Yd= Y - T
Y = C + I + G + (X - M)
Y = 2000 + 0,75*(Y-100-0,25Y)+2000+4600+(2550-410-0,3Y)
Y-0,75Y+0,1875Y+0,3Y = 8600-75+2550-410
0,7375Y = 10 665
Ye = 14 461 (the level of equilibrium national income)